OpenStudy (anonymous):

2+1/2+1/2_ _ _ _ to infinite times. What will be the value of the series?

OpenStudy (experimentx):

what are the few other terms??

OpenStudy (anonymous):

@ashishthomas7 , your 3rd term is incorrect. Check that :)

OpenStudy (anonymous):

thats it 2+1/2+1/2+1/2+........to infinte terms

OpenStudy (anonymous):

it would be infinity then no?

OpenStudy (experimentx):

2 + (1/2 + 1/2 + 1/2 + ..................+1/2+...........) ?? that would be infinity

OpenStudy (anonymous):

nope..its not the answer that i have in my book..

OpenStudy (anonymous):

whats the answer given?

OpenStudy (anonymous):

1+2^1/2

OpenStudy (anonymous):

Well, Use \[2 + \frac{1}{2} (1+1+1.............\infty)\] \[2 + \frac{1}{2}( \infty)\] \[2+ \infty\] \[\infty\] This should be the answer unless you have mistyped the question or I have made some mistake

OpenStudy (anonymous):

its not infinity.

OpenStudy (anonymous):

i meant infinte times

OpenStudy (anonymous):

i have typed the question correctly

OpenStudy (anonymous):

1+1+1+ .......(infinite times ) = infinte

OpenStudy (experimentx):

it should have been like this \[ 2(1+1/2-1/8+1/16-5/128+7/256 + .... ) \] try to add few terms, obviously you will get the answer is invalid

OpenStudy (anonymous):

@ashishthomas7 , can you take a screenshot of the question and attach it here :)

OpenStudy (anonymous):

\[\sqrt{6+\sqrt{6+\sqrt{6+\infty}}}\]its a type of this question actually edited the stuff..

OpenStudy (anonymous):

Yes \[y =\sqrt{6+\sqrt{6+\sqrt{6+\infty}}}\] \[y^2 = 6+ y\]

OpenStudy (anonymous):

That's all :D

OpenStudy (anonymous):

You get a quadratic equation. Now solve and get your y value

OpenStudy (anonymous):

@ashishthomas7

OpenStudy (anonymous):

well now can u do that in this series 2+\[2+1/2+1/2+.........\infty\]

OpenStudy (anonymous):

How ??

OpenStudy (anonymous):

i think we hav to do like this :\[2+1/2+1/2+.....\infty\] =\[2+1/X=X\]

OpenStudy (anonymous):

@ashishthomas7 , Check again what you wrote .

OpenStudy (anonymous):

i dont have the book :( but the question iscertainly like this at a certain level u will get it \[2x+1=x ^{2}\]

OpenStudy (anonymous):

do you mean a repeated fraction?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

According to you, \[2+ \frac{1}{2+1/2 +1/2+1/2 +.......} = 2+1/2 +1/2 +...\] How is this possible??

OpenStudy (anonymous):

notlike that

OpenStudy (anonymous):

(2+1/(2+1/(2+1/(2+....

OpenStudy (anonymous):

That's what you wrote @ashishthomas7 @eigenschmeigen , his question is different

OpenStudy (anonymous):

\[2+1/2+1/2+1/2+..........x terms\] now guys try

OpenStudy (anonymous):

the question is not what yopu think it should be \[2+1\div(2+1\div(2+......))\]

OpenStudy (anonymous):

Yes Exactly. That's why I am asking for some source / screenshot of the question

OpenStudy (anonymous):

that means it turns out to be \[x ^{2}=2x+1\] => x=\[1+\sqrt{2}\]

OpenStudy (anonymous):

yup thats the answer.

OpenStudy (anonymous):

@ashishthomas7 , please check your question before writing :)

OpenStudy (anonymous):

I meant next time

OpenStudy (anonymous):

@shivam_bhalla sorry :)

OpenStudy (anonymous):

It's Ok :D

OpenStudy (shubhamsrg):

is it  ?